I plan on retiring in 35 years. I hope to have $750,000 in my savings account when I retire. If I found a savings account that earns 6.4% interest compound monthly how much should I put in my savings account now to reach my goal?
Answer by Boreal(15235)

750000=d((1+.064/12)^420-1)/.064/12
multiply through by 0.064/12
4000=d(9.338-1)=8.338, but this is rounded. Keep it unrounded on the calculator
d=$479.96